Week of 200120

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Julia (WLCG), Maarten (ALICE, WLCG), Kate (WLCG, chair), Petr (ATLAS), Julien (storage), Olga (computing), Borja (monitoring)
  • remote: Elena (CNAF), Onno (NL-T1), David B (IN2P3), Dave M (FNAL), Christoph (CMS), Di (TRIUMF), Sang Un (KISTI), Andrea V (LHCb)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Work in progress with Grand Unification ADCDPA-235 - allow our WFMS to make more optimal decisions for Analysis vs. Production
    • 3 CERN FTS with failing transfers INC2274140
      • No FTS server has updated the transfer status the last 900 seconds. Probably stalled
    • Few issues with slow queries in Rucio (related to new features from October release)
    • Slow release of storage space for deleted files at BNL GGUS:144845
    • Starting with full Run2 reprocessing campaign using data carousel ADCDPA-236
    • Storage permissions INC2277811

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Heavy staging at sites will start again during the week because of stripping campaign

Sites / Services round table:

  • ASGC:
  • BNL:
  • CNAF: On January 23 there is a scheduled downtime (from 00 to 20.00 CET) in order to upgrade TSM tape server. Recalls will be put in pending state, while the disk buffer will remain always available. GOCDB Link
  • EGI:
  • FNAL: Planning downtime for tape dCache instance on 1/29 to upgrade to v5.2, then disk dCache following about two weeks later.
  • IN2P3: NTR
  • JINR: NTR
  • KISTI: NTR
  • KIT:
    • Resolved issues with updating CRLs (GGUS:144814)
    • Networking issues in KIT between Friday evening and Saturday morning did not affect GridKa operations, as far as we know.
    • Declared downtime at-risk from Thursday 1 pm till 5 pm CET, where we want to update the firmware of a troublesome storage system. The redundancy should make this transparent, but it may also break the storage availability for everyone, if things go badly.
  • NDGF:
  • NL-T1: NTR
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL:
  • TRIUMF: NTR

  • CERN computing services: OTG:0054182 - Some batch nodes not available due to network switch outage.
  • CERN storage services:
    • Due to planned router intervention by Network Team in DC (Mon 20th 15:00-17:30, Wed 22th 14:30-16:30, access to some EOS files can be affected, Should be mostly transparent since only a few disk servers are affected (OTGs: OTG:0054176, OTG:0054177)
    • The announced interventions on the CASTOR Stager databases have been published in the Service Status Board and are planned as follows, if no objections arise:
  • CERN databases:
  • GGUS:
    • On Jan 13 the support privileges of expert users were accidentally lost.
      • Such experts are allowed to open multiple tickets in one go.
      • It was deemed too painful to restore a DB backup just for that.
      • Instead, the lists are being repopulated from VOMS groups/roles per experiment.
  • Monitoring:
    • Final availability reports for Dec 2019 sent around
  • MW Officer:
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r23 < r22 < r21 < r20 < r19 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r23 - 2020-01-27 - GiuseppeLoPresti
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback